OpenStudy (egbeach):

Express the complex number in trigonometric form. 3 - 3i

OpenStudy (loser66):

On it, what is x? what is y?

OpenStudy (egbeach):

x=3 y=3i right?

OpenStudy (egbeach):

@Loser66

OpenStudy (loser66):

x =3, y =3 , not 3i , ok?

OpenStudy (loser66):

now \(x^2 + y^2 =r^2\) , hence r =??